# Best Agent Identification for General Game Playing

### v1

@@ 3.3.1 GapE
The Gap-based Exploration (GapE) algorithm (Gabillon et al., 2011) is one of the first approaches designed specifically for the multi-bandit best arm identification problem. This algorithm favours arms with an expected reward close to the best performing arm (i.e., those with the smallest “gap”). The authors of this approach also define a variant called GapE-V, which additionally considers the variance in rewards for each arm.
@@
While on the surface this approach seems well suited for our intended application, the specific characteristics of our general game playing domains highlight some potential issues. Prior observations and experience with these general game frameworks has demonstrated that, while there are certain games where agents will frequently achieve different outcomes between repeated trials, a significant portion of games also have a very low variability in agent performance. In these cases, there are often several agents that either always win or always lose the game (i.e., reward distributions for game-agent pairs are asymmetric and centred towards the extremities). This means that there will be multiple arms for a given bandit with a near-zero gap between them and the best performing arm.
@@
For our proposed domain, we care primarily about minimising the average simple regret $r(n)$ of our selected arms. As such, if there are several best or near-best arms for a given game, then the decision of which one we select will have a negligible impact. Because GapE has been designed primarily for minimising the average probability of error $e(n)$, the approach instead focuses a large number of arm pulls on these equally high performing arms (i.e., the algorithm attempts to identify the one true “best” arm, rather than being directly concerned with reducing the simple regret of the bandit). This difference often leads to a sub-optimal arm selection strategy from GapE when attempting to minimise the average simple regret across all bandits.
3.3.2 UCB-E
@@
The UCB-E algorihtm (Audibert et al., 2010) is a highly exploring variant of the Upper Confidence Bound (UCB) selection policy (Auer, 2003). One of the most common algorithms for standard (i.e., cumulative reward) multi-arm bandit problems is the UCB1 algorithm (Auer et al., 2002), which is able to effectively balance exploration and exploitation. However, best arm identification can be considered a pure exploration problem, as the rewards obtained from each arm during runtime do not affect our performance score (only the final arm pull does). UCB-E attempts to increase the degree of exploration performed by UCB, based on a defined exploration value $a$. If $a=2\cdot\log(n)$, where $n$ is the current number of rounds, then UCB-E is equivalent to UCB1, with the value for $a$ being increased to incentivise more exploration. The effectiveness of UCB-E is highly dependant on the value chosen for $a$, for which the optimal value is often unknown and cannot be easily estimated from past observations. While originally designed for best arm identification in a single bandit case, UCB-E can be generalised to the multi-bandit environment by selecting bandits uniformly and then pulling arms within each bandit using the UCB-E algorithm (Gabillon et al., 2011).
3.3.3 Successive Rejects
@@
The Successive Rejects algorithm (Audibert et al., 2010) repeatedly removes the worst performing arm over multiple rounds of selection until just a single arm remains, a process commonly referred to as “action elimination”(Jamieson and Nowak, 2014). However, in order for this and other action elimination algorithms to be used to their fullest effectiveness, the total number of arm pulls must be known in advance. This lack of a “stop anytime” functionality is a significant downside compared to previous approaches, particularly for our intended use case of general game playing. The amount of time needed to complete a single game trial can be highly variable, meaning that we cannot easily estimate the exact number of trials that can be run within an allotted time. The generalisation of this approach to multiple bandits is also non-optimal, requiring the number of rounds (i.e., arm pulls) to be equally split among all bandits at the start of the process, rather than to bandits with the highest potential benefit during runtime.
3.3.4 Sequential Halving
@@
The original Sequential Halving algorithm (Karnin et al., 2013) works in a manner similar to that of the Successive Rejects algorithm, where the worst performing arms of each bandit are removed at the end of each of multiple rounds. However, rather than removing a single arm at the end of each round, half of the remaining arms are removed instead. Sequential Halving also suffers from the same core limitation as Successive Rejects, namely being the requirement that the number of available arm pulls must be known in advance.
@@
Anytime Sequential Halving (Sagers et al., 2025) provides a variant approach that avoids this limitation. This variant performs an initial run of the Sequential Halving algorithm, with a limited budget of arm pulls equal the minimum amount required for a single complete run. Once complete, this “minimal pull run” process is repeated, with subsequent reward estimates for each arm incorporating all prior pull results from previous runs. While this Anytime Sequential Halving variant offers the desired stop anytime functionality, it still suffers from being unable to generalise effectively to the multi-bandit domain. Similar to Successive Rejects, arm pulls must be divided evenly across all bandits, resulting in an equal number of trials for each game.
3.4 Wilson Score
The Wilson score interval is a binomial proportion confidence interval for estimating the probability of success based on prior Bernoulli (aka. binomial) trial outcomes (Wilson, 1927). The Wilson score interval has previously been shown to perform particularly well compared to other confidence estimates when the number of samples is small, or when the probability of success is close to the extremes (i.e., 0 or 1) (Agresti and Coull, 1998; Wallis, 2013). The Wilson score interval for a specified confidence level $(1-\alpha)$ is defined in formula (1), additionally parameterised by the sample probability of success $\hat{p}$ and number of samples $n$. Note, for the sake of our later algorithms, we assume that if the value for $n$ is zero then the returned Wilson score interval is always $(0,1)$.
$Wilson(\hat{p},n,\alpha)\equiv(w^{-},w^{+})=\frac{\hat{p}+\frac{z^{2}_{\alpha/2}}{2n}\pm z_{\alpha/2}\sqrt{\frac{\hat{p}(1-\hat{p})}{n}+\frac{z^{2}_{\alpha/2}}{4n^{2}}}}{1+\frac{z^{2}_{\alpha/2}}{n}}$ (1)
While the Wilson score interval was originally intended only for Bernoulli trials with an outcome of either 0 or 1, it can still be applied in practice to non-Bernoulli trials as long as the outcome is always within the required $[0,1]$ bounds. In this case, the value $\hat{p}$ is replaced with the average result of all trials. The fact that Wilson score intervals are known to work particularly well with highly skewed distributions (Wallis, 2013), suggests that this generalised interpretation may provide a suitable confidence estimate for agent win-rate in our general game playing domain.
4 Optimistic-WS
@@
Based on these criteria, we propose our Optimistic-WS algorithm to tackle the best agent identification problem for general game playing. Algorithm 1 provides the pseudocode for this approach, utilising previously defined variables and terminology. A more descriptive explanation of Optimistic-WS operations and the motivation behind its design is provided below.
Algorithm 1 Pseudocode of the Optimistic-WS algorithm
number of rounds $n$, exploration value $c$
@@
              Compute $(w^{-}_{mk},w^{+}_{mk})=Wilson(\hat{\mu}_{mk}(t-1),T_{mk}(t-1),\alpha)$
@@
The general philosophy behind our proposed Optimistic-WS algorithm is to take the same “optimism in the face of uncertainty” perspective used for most upper confidence bound selection policies. In this case, we selected the Wilson score interval to estimate the upper and lower confidence bounds $[w^{-}_{mk},w^{+}_{mk}]$ of the expected reward of each bandit-arm pair $(m,k)$. The confidence bounds for a given arm are determined based on its expected return $\hat{\mu}_{mk}$, as well as the number of times it has previously been pulled $T_{mk}$ and the confidence value $\alpha$. Higher values for $T_{mk}$ and $\alpha$ will shrink the confidence interval (vice versa). Due to the fact that $Wilson(\hat{p},n,\alpha)$ always returns the maximal confidence interval $(0,1)$ when $n$ is zero, any arm that has not yet been pulled at least once (i.e., $T_{mk}=0$) will always be the preferred selection choice.
While a fixed confidence value ($\alpha$) could be used for all rounds, we instead chose to apply an adaptive confidence value that decreases as the round number ($t$) increases. This reduction ensures that our arm selection process will continue to explore across all bandits even for a very high number of rounds, and guarantees that our proposed Optimistic-WS algorithm will converge towards the optimal result as the number of rounds increases. Without this, it is possible that the gap between the confidence values of the best and second best arms in a bandit becomes large enough that no arms of this bandit are ever pulled again. In practice, the rate at which our confidence value decreases can be adjusted using a specific exploration value ($c$), with the total number of arms $K_{m}$ used as a stabilising value across different problem domains.
The potential regret change $\Delta_{mk}$ for each bandit-arm pair $(m,k)$ is then calculated by comparing its Wilson score confidence bounds $[w^{-}_{mk},w^{+}_{mk}]$ against the expected reward for the current best predicted arm $\hat{\mu}_{m}^{\star}$. If an arm $k$ is not the current best predicted arm $\hat{k}_{m}^{\star}$ for a bandit $m$, then its potential regret change $\Delta_{mk}$ is equal to the difference between its upper confidence bound $w^{+}_{mk}$ and the expected reward of the best performing arm $\hat{\mu}_{m}^{\star}$. This represents the potential increase in regret that would be obtained by selecting arm $k$ for bandit $m$, if its true expected reward were equal to its upper confidence bound $w^{+}_{mk}$. Conversely, if the arm $k$ is the best predicted arm $\hat{k}_{m}^{\star}$ for a bandit $m$, then its potential regret change $\Delta_{mk}$ is equal to the difference between the expected reward of the best performing arm $\hat{\mu}_{m}^{\star}$ (which in this case will be identical to its own expected reward $\hat{\mu}_{mk}$) and its lower confidence bound $w^{-}_{mk}$. This represents the potential decrease in regret that would be obtained by selecting arm $k$ for bandit $m$, if its true expected reward was equal to its lower confidence bound $w^{-}_{mk}$. In summary, we apply an optimistic perspective on all non-best arms by assuming their upper confidence bound is true, and a pessimistic perspective on all best arms by assuming their lower confidence bound is true.
@@
The bandit-arm pair with the largest $\Delta_{mk}$ value is then pulled to obtain our new reward sample $X_{mk}$. This reward sample is used to update the arm’s expected reward estimate $\hat{\mu}_{mk}$ and increment its pulls counter $T_{mk}$. Once this process has been completed the round is over, and the next round can begin. The current best predicted arm for any bandit $J_{m}(t)$ can be returned at the end of any round $t$, providing an intermediate measure of algorithm performance. New bandits or arms can also be added at the start of any round by updating the values for $M$ and $K_{m}$, as well as initialising all values for $T_{mk}(0\dots t)$ and $\hat{\mu}_{mk}(0\dots t)$ to 0.
4.1 Hyperparameters
4.1.1 Number of rounds

### v2

@@ III-C1 GapE
The Gap-based Exploration (GapE) algorithm [undefq] is one of the first approaches designed specifically for the multi-bandit best arm identification problem. This algorithm favours arms with an expected reward close to the best performing arm (i.e., those with the smallest “gap”). The authors of this approach also define a variant called GapE-V, which additionally considers the variance in rewards for each arm.
@@
While on the surface this approach seems well suited for our intended application, the specific characteristics of our general game playing domain highlight some potential issues. Prior observations and experience with these general game frameworks has demonstrated that, while there are certain games where agents frequently achieve different outcomes between repeated trials, a significant portion of games also have a very low variability in agent performance. In these cases, there are several agents that either always win or always lose the game (i.e., reward distributions for game-agent pairs are asymmetric and centred towards the extremities, see Figures 1 and 2). This means that there are often multiple arms for a given bandit with a near-zero gap between them and the best performing arm.
@@
For our proposed domain, we care primarily about minimising the average simple regret $r(n)$ of our selected arms. As such, if there are several best or near-best arms for a given game, then the decision of which one we select has a negligible impact. GapE appears to have been designed primarily for minimising the average probability of error $e(n)$, and so it instead focuses a large number of arm pulls on these equally high performing arms (i.e., attempts to identify the “one true” best arm, rather than being directly concerned with reducing the simple regret of the bandit). This difference often leads to a sub-optimal arm selection strategy from GapE when attempting to minimise the average simple regret across all bandits.
III-C2 UCB-E
@@
The UCB-E algorithm [undefd] is a highly exploring variant of the Upper Confidence Bound (UCB) selection policy [undefe]. One of the most common algorithms for standard (i.e., cumulative reward) multi-arm bandit problems is the UCB1 algorithm [undeff], which is able to effectively balance exploration and exploitation. However, best arm identification can be considered a pure exploration problem, as the rewards obtained from each arm during runtime do not affect our performance score (only the final arm pull does). UCB-E attempts to increase the degree of exploration performed by UCB, based on a defined exploration value $a$. If $a=2\cdot\log(n)$, where $n$ is the current number of rounds, then UCB-E is equivalent to UCB1, with the value for $a$ being increased to incentivise more exploration. The effectiveness of UCB-E is highly dependant on the value chosen for $a$, for which the optimal value is often unknown and cannot be easily estimated from past observations. While originally designed for best arm identification in a single bandit case, UCB-E can be generalised to the multi-bandit environment by selecting bandits uniformly and then pulling arms within each bandit using the UCB-E algorithm [undefq].
III-C3 Successive Rejects
@@
The Successive Rejects algorithm [undefd] repeatedly removes the worst performing arm over multiple rounds of selection until just a single arm remains, a process commonly referred to as “action elimination” [undefy]. However, in order for this and other action elimination algorithms to be used to their fullest effectiveness, the total number of arm pulls must be known in advance. This lack of a “stop anytime” functionality is a significant downside compared to previous approaches, particularly for our intended use case of general game playing. The amount of time needed to complete a single game trial can be highly variable, meaning that we cannot easily estimate the exact number of trials that can be run within an allotted time. The generalisation of this approach to multiple bandits is also non-optimal, requiring the number of rounds (i.e., arm pulls) to be equally split among all bandits at the start of the process, rather than to bandits with the highest potential benefit during runtime.
III-C4 Sequential Halving
@@
The original Sequential Halving algorithm [undefaa] works in a manner similar to that of the Successive Rejects algorithm, where the worst performing arms of each bandit are removed at the end of each of multiple rounds. However, rather than removing a single arm at the end of each round, half of the remaining arms are removed instead. Sequential Halving also suffers from the same core limitation as Successive Rejects, namely being the requirement that the number of available arm pulls must be known in advance.
@@
Anytime Sequential Halving [undefal] provides a variant approach that avoids this limitation. This variant performs an initial run of the Sequential Halving algorithm, with a limited budget of arm pulls equal the minimum amount required for a single complete run. Once complete, this “minimal pull run” process is repeated, with subsequent reward estimates for each arm incorporating all prior pull results from previous runs. While this Anytime Sequential Halving variant offers the desired stop anytime functionality, it still suffers from being unable to generalise effectively to the multi-bandit domain. Similar to Successive Rejects, arm pulls must be divided evenly across all bandits, resulting in an equal number of trials for each game.
III-D Confidence Intervals
Our proposed RCP algorithm (introduced in the next section) relies on a provided confidence interval to estimate the upper and lower reward bounds for a specific arm (based on prior results). While the RCP algorithm can utilise any given upper/lower bound calculation, it should ideally employ a confidence interval that accurately represents the underlying reward distribution. We have therefore chosen to test a variety of suitable confidence interval estimation formulas, evaluating which is most suitable for our intended domain of general game playing.
The majority of confidence intervals described in this section were originally intended for modelling Binomial distributions, where each event has an outcome of either 0 or 1. However, a generalised version of these formulas can still be applied in approximation to non-Bernoulli trials, so long as the outcome is always within the required $[0,1]$ bounds [undefw]. In this case, the proportion of successes ($\hat{p}$) is replaced with the average result of all trials. This generalisation is supported by the Bhatia-Davis inequality, $\sigma^{2}\leq\mu(1-\mu)$, which guarantees the true variance is always less than or equal to that assumed by the binomial confidence interval [undefh]. This allows for binomial confidence intervals to be used as a conservative confidence interval estimation in non-Bernoulli cases.
The decision to focus primarily on binomial confidence intervals over more general distribution options is based on the knowledge that the majority of general game outcomes are either a win or a loss. All games used in the GVGAI framework are single-player, meaning that every trial result is either a 0 (loss) or 1 (win). However, the Ludii general game playing system also contains multi-player games, where there is the possibility of a 0.5 outcome for a draw and fractional value outcomes for games with multiple opponents. While this means that the results of Ludii games are not strictly binomial in every case, the fact that the vast majority of game outcomes still end in either a win or loss means that they still exhibit highly binomial distribution patterns in practice (resulting in the need for a generalised binomial distribution interpretation). For alternative problem domains, it may be beneficial to experiment with alternative confidence interval estimations more suited to the underlying result distribution.
III-D1 Wilson Score
The Wilson score interval is an approximate binomial proportion confidence interval obtained by inverting a score test for the Bernoulli parameter [undefaaa]. Unlike the traditional Wald interval, the Wilson interval avoids relying solely on the asymptotic normality of the sample proportion and instead incorporates an adjustment based on the test statistic. This leads to substantially improved coverage behaviour, particularly when sample sizes are small or when the observed probability of success lies near the extremes (i.e., 0 or 1) [undef, undefay]. The Wilson score interval for a specified confidence level $(1-\alpha)$ is defined in formula (1), additionally parameterised by the proportion of successes $\hat{p}$ and sample size $n$.
$\footnotesize Wilson(\hat{p},n,\alpha)\equiv(b^{-},b^{+})=\frac{\hat{p}+\frac{z^{2}_{\alpha/2}}{2n}\pm z_{\alpha/2}\sqrt{\frac{\hat{p}(1-\hat{p})}{n}+\frac{z^{2}_{\alpha/2}}{4n^{2}}}}{1+\frac{z^{2}_{\alpha/2}}{n}}$ (1)
III-D2 Jeffreys
The Jeffreys interval is a Bayesian binomial proportion confidence interval, derived from the Jeffreys prior for the Bernoulli parameter which corresponds to a $Beta(\frac{1}{2},\frac{1}{2})$ distribution [undefz]. Similar to the Wilson score interval, Jeffreys interval has been shown to perform well in small sample and unbalanced heavy distributions [undefj]. For a specified confidence level $(1-\alpha)$ the Jeffreys interval is obtained by taking the lower and upper $\alpha/2$ and $1-\alpha/2$ quantiles of the posterior Beta distribution, see formula (2), additionally parameterised by the number of successes $\hat{x}$ and sample size $n$.
$Jeffreys(\hat{x},n,\alpha)\equiv\Big($ $\,Q_{\alpha/2}\Big(\hat{x}+\frac{1}{2},\,n-\hat{x}+\frac{1}{2}\Big),$ (2)
$\,Q_{1-\alpha/2}\Big(\hat{x}+\frac{1}{2},\,n-\hat{x}+\frac{1}{2}\Big)\Big)$
III-D3 Clopper–Pearson
The Clopper–Pearson interval is an exact binomial proportion confidence interval derived by inverting the cumulative distribution function of the binomial distribution [undefm]. Unlike approximate methods, it provides guaranteed frequentist coverage for all parameter values, although this guarantee typically results in intervals that are conservative, particularly when the sample size is small or when the observed proportion is near 0 or 1 [undefj]. For a specified confidence level $(1-\alpha)$ the lower and upper bounds of the Clopper–Pearson interval are obtained from the $\alpha/2$ and $1-\alpha/2$ quantiles of Beta distributions, see formula (3), additionally parameterised by the number of successes $\hat{x}$ and sample size $n$.
$ClopperPearson(\hat{x},n,\alpha)\equiv\Big($ $\,Q_{\alpha/2}\Big(\hat{x},\,n-\hat{x}+1\Big),$ (3)
$)\displaystyle\,Q_{1-\alpha/2}\Big(\hat{x}+1,\,n-\hat{x}\Big$$\Big)$
III-D4 Agresti–Coull
The Agresti–Coull interval is an approximate binomial proportion confidence interval that improves upon the traditional Wald interval by adding $z_{\alpha/2}^{2}$ pseudo‑observations, split equally between wins and losses, resulting in an adjusted sample size and proportion [undef]. This modification yields intervals with markedly better coverage properties, particularly for small sample sizes or when the observed proportion is close to 0 or 1 [undefj]. For a specified confidence level $(1-\alpha)$, the Agresti–Coull interval is constructed using the adjusted sample size $\tilde{n}=n+z_{\alpha/2}^{2}$ and proportion of successes $\tilde{p}=(\hat{x}+z_{\alpha/2}^{2}/{2})/\tilde{n}$, from which the adjusted confidence interval is computed, see formula (4).
$\footnotesize AgrestiCoull(\hat{p},n,\alpha)\equiv(b^{-},b^{+})=\tilde{p}\pm z_{\alpha/2}\sqrt{\frac{\tilde{p}(1-\tilde{p})}{\tilde{n}}}$ (4)
III-D5 Kullback–Leibler Upper Confidence Bound (KL-UCB)
The KL‑UCB interval constructs confidence bounds for Bernoulli rewards by inverting a constraint on the Kullback–Leibler divergence, rather than using distribution‑free inequalities [undefs]. Given the proportion of successes $\hat{p}$, the number of arm samples $n$, and the total number of samples across all arms $t$, the upper and lower bounds are defined as the solutions $b^{+}\!\geq\!\hat{p}$ and $b^{-}\!\leq\!\hat{p}$ to the implicit equations $n\cdot D_{\mathrm{KL}}(\hat{p}\,\|\,q)\leq log(t)+c\cdot log(log(t))$, where $D_{\mathrm{KL}}$ is the Bernoulli KL divergence and $c$ is an exploration constant controlling confidence. Compared with more standard Hoeffding‑type intervals, the KL‑UCB interval yields tighter bounds near the extremes and improved exploration–exploitation trade‑offs [undefl].
III-D6 Probably Approximately Correct (PAC)
The Probably Approximately Correct (PAC) interval is a binomial proportion confidence interval derived from Hoeffding‑style guarantees, ensuring that the empirical success rate remains within a fixed deviation of the true parameter with high probability [undefaw]. In our setting, we adopt an adaptive confidence allocation that spreads the total error budget $\delta$ across all bandits, arms, and number of samples using a $\sum_{t\geq 1}1/{t^{2}}$ normalisation. Concretely, for an observed proportion of successes $\hat{p}$, arm sample count $n$, number of bandits $M$, number of arms $K$, and total samples $t$, we set:
$\footnotesize\alpha_{\mathrm{adapt}}=\frac{6\,\delta}{\pi^{2}MKt^{2}}\qquad\beta_{\mathrm{adapt}}=\sqrt{\frac{\log\!\left(2/\alpha_{\mathrm{adapt}}\right)}{2\max(1,n)}}$ (5)
We then define the confidence interval as:
$\footnotesize PAC(\hat{p},n,M,K,t,\delta)\equiv\left(\max\{0,\,\hat{p}-\beta_{\mathrm{adapt}}\},\;\min\{1,\,\hat{p}+\beta_{\mathrm{adapt}}\}\right)$ (6)
This adaptation yields a union‑bound‑controlled guarantee over all bandit arms, with the confidence interval width shrinking at rate $1/\sqrt{n}$, while also growing via the $t^{2}$ weighting. This ensures that the upper/lower confidence bounds for each arm continually increases/decreases (respectively) so long as the arm in question is not pulled. When utilising this confidence interval, our RCP method achieves a fixed-confidence guarantee (proof provided in Appendix D).
IV Regret Change Potential (RCP)
@@
Based on these criteria, we propose our RCP algorithm to tackle the best agent identification problem for general game playing. Algorithm 1 provides the pseudocode for this approach, utilising previously defined variables and terminology, when employing the Wilson Score confidence interval. Similar versions of this algorithm can be created for other confidence intervals, by replacing the $Wilson(\hat{p},n,\alpha)$ call with the desired confidence interval function (e.g., $Jeffreys(\hat{x},n,\alpha)$ or $AgrestiCoull(\hat{p},n,\alpha)$).
Algorithm 1 Pseudocode of the RCP algorithm (when using the Wilson Score interval)
number of rounds $n$, alpha value $\alpha$
@@
     Compute $(b^{-}_{mk},b^{+}_{mk})=Wilson(\hat{\mu}_{mk}(t-1),T_{mk}(t-1),\alpha)$
@@
[…]
@@
The bandit-arm pair with the largest $\Delta_{mk}$ value is pulled to obtain our new reward sample $X_{mk}$, which is then used to update the parameters of our chosen confidence bound (e.g., for the Wilson score interval, this is the arm’s expected reward estimate $\hat{\mu}_{mk}$ and pulls counter $T_{mk}$). Once this process has been completed the round is over, and the next round can begin. The current best predicted arm for any bandit $J_{m}(t)$ can be returned at the end of any round $t$, providing an intermediate measure of algorithm performance. New bandits or arms can also be added at the start of any round by updating the values for $M$ and $K_{m}$, as well as initialising all values for $T_{mk}(0\dots t)$ and $\hat{\mu}_{mk}(0\dots t)$ to 0. Once the specified number of rounds is completed, for each bandit ($\forall_{m}\in\{1\dots M\}$) we return the arm with the highest expected reward ($J_{m}(n)=\operatorname*{argmax}_{k}\hat{\mu}_{mk}$).
IV-A Parameters
IV-A1 Number of rounds